Shift report 20020110 16:00-24:00

From: Apache (apache@pii3.brahms.bnl.gov)
Date: Thu Jan 10 2002 - 23:46:28 EST

  • Next message: Hironori Ito: "test test test."

    ------------------------------------------------------------
    Date: Thu Jan 10 23:46:27 2002, Shift: 16:00-24:00
    Supervisor: IGB, with: KH
    Log book#12 pages 42 - 43
    
    Shift Summary:
    ------------------------------------------------------------
    Machine Status:
    Bad beam from 1830-2100, beam had large emmitance and had to  be dumped.
    No beam 2100-2200.
    Good beam from 2200.
    
    Achievements:
    Tricked Kris Hagel into taking the first half of the shift.
    Have taken (so far) roughly 100K MRS, 35K FS with 
    MRS at 45, 4B and FS at 12 1/4 field.
    
    Detector problems:
    None noticed.
    ------------------------------------------------------------
    



    This archive was generated by hypermail 2b30 : Thu Jan 10 2002 - 23:47:22 EST